### v1

# A Survey on SAR ship classification using Deep Learning

Abstract
Deep learning (DL) has emerged as a powerful tool for Synthetic Aperture Radar (SAR) ship classification. This survey comprehensively analyzes the diverse DL techniques employed in this domain. We identify critical trends and challenges, highlighting the importance of integrating handcrafted features, utilizing public datasets, data augmentation, fine-tuning, explainability techniques, and fostering interdisciplinary collaborations to improve DL model performance. This survey establishes a first-of-its-kind taxonomy for categorizing relevant research based on DL models, handcrafted feature use, SAR attribute utilization, and the impact of fine-tuning. We discuss the methodologies used in SAR ship classification tasks and the impact of different techniques. Finally, the survey explores potential avenues for future research, including addressing data scarcity, exploring novel DL architectures, incorporating interpretability techniques, and establishing standardized performance metrics. By addressing these challenges and leveraging advancements in DL, researchers can contribute to developing more accurate and efficient ship classification systems, ultimately enhancing maritime surveillance and related applications.
Index Terms:
SAR ship classification, Deep learning, Synthetic Aperture Radar
I Introduction
Our planet is covered by approximately 71% water, which facilitates over 80% of global trade. This vast commercial activity necessitates maritime surveillance, which is a complex research field due to its extensive coverage and diverse monitoring requirements. These requirements encompass irregular migration, piracy, fisheries management, and traffic monitoring. Maritime surveillance plays a crucial role in promoting environmental sustainability by providing real-time data and insights to tackle pollution, reduce emissions, conserve marine ecosystems, and ensure the long-term health of our oceans.
Many common applications of maritime monitoring exploit SAR and Automatic Identification System (AIS) data. SAR data can be collected from satellites and airplanes, with the ability to acquire information about objects on the sea surface, such as ships. The benefits of SAR are that it can operate in any light and weather conditions, it can provide images of a vast area and revisit the same area repeatedly. AIS data are transmitted from ships, received and broadcasted by ground or satellite-based stations, and contain information about the vessel, such as its unique ID, position, speed, course, and other relevant details. These data are frequently transmitted for applications such as maritime traffic management, collision avoidance, search and rescue. AIS and SAR data can be combined to create datasets that enhance maritime monitoring applications.
A critical feature of SAR technology is polarization, which refers to the orientation of the electromagnetic waves used in the radar signals. SAR systems can operate in various polarization modes, such as single-polarization (like VV or HH), dual-polarization (such as VV and VH, or HH and HV), and fully-polarized modes (VV, VH, HH, HV). Here, ”V” stands for vertical polarization and ”H” for horizontal polarization, with the first letter indicating the transmitted polarization and the second the received polarization. These different modes provide varying levels of detail and information about the target, influencing the type of analysis and the performance of specific deep learning algorithms. Different acquisition modes, such as single-look or multi-look, further influence image characteristics. Understanding these fundamental aspects of SAR imagery is essential for effectively utilizing the information contained within these datasets.
Due to coverage, availability and benefits of SAR data, researchers were able to use it for several applications such as Oil Spill Detection [1, 2, 3], Ship Detection [4, 5, 6], Ship velocity estimation [7, 8, 9, 10], and Ship classification [11, 12, 13, 14]. In SAR images, the area represented by each pixel varies depending on the specific sensing platform. For example, the Sentinel-1 C-band SAR satellite constellation provides images of Earth’s surface with a spatial resolution of approximately 20 meters. This limited resolution can render vessels as only a few pixels, making object identification challenging without additional information. To address this limitation and achieve an accurate target labeling, when possible, SAR-derived objects are integrated with AIS data [15].
The SAR data are matched with AIS data using several techniques [16] such as time-based matching, range-based matching and feature-based matching. The SAR images are cropped so that each image chip represents a single detected ship, then labelled according to the matched AIS data and, finally, different Machine Learning models are trained to detect and classify ships from unseen SAR images. The output is usually a set of bounding boxes in the large SAR image indicating the presence of the detected ships and their associated categories. This approach to classification has shown good results using machine learning models, and also deep learning (DL) models are being explored to the same purpose.
Ships exhibit a remarkable diversity in types. For example, the FUSAR dataset [17] encompasses 15 primary ship categories and 98 subcategories. While this extensive dataset proves valuable, ship classification remains a challenging task due to several factors. The limited resolution of SAR images often obscures the unique characteristics of different ships. Additionally, their single-channel nature, as opposed to the multichannel images commonly available from optical sensors, limits the use of color-based features, further complicating object classification. Moreover, the prevalence of cargo ships, the most widely used type, introduces a significant class imbalance, making cargo ships easier to be classified compared to less represented vessels. Machine learning models often struggle with ship classification due to the absence of a clear understanding of which features are most relevant for accurate classification.
Conversely, DL approaches offer a promising solution for ship classification challenges: they can autonomously extract relevant features from satellite imagery, thus aiding vessel identification. However, the effectiveness of DL models heavily relies on crucial decisions regarding data selection, augmentation techniques, learning algorithms, and model architectures. This survey provides a comprehensive review of various techniques employed to train DL models for ship classification.
While existing surveys focus on SAR ship detection using DL [5, 4], a dedicated survey on SAR ship classification using DL is currently absent. This survey addresses a critical gap within the scientific community by providing a comprehensive overview of the state of SAR ship classification using DL. It aims to bridge this gap by summarizing the existing literature, elucidating key methodologies, challenges, and advancements in the field. Additionally, the survey offers insights into past developments, current trends, and future directions in SAR ship classification, thus serving as a valuable resource for researchers and practitioners in this specialized domain. The following chapters outline the structure of this survey: Chapter 2 details the methodology employed for performing survey; Chapter 3 presents the survey’s findings resulting in a DL taxonomy; Chapter 4 reviews relevant literature; Chapter 5 explores key insights from the literature review; Chapter 6 access the design consideration for DL SAR Ship classification; Chapter 7 discusses the key issues; Chapter 8 outlines the future; and Chapter 9 concludes the paper.
II Survey Methodology
II-A Research Questions
The aim of this survey is to find and highlight techniques that are being used in improving ship classification through DL. The research questions are focused on finding the concerning literature:
    RQ1: How has the use of specific DL techniques evolved over time in the field of SAR ship classification?
    RQ2: What are the performance characteristics (e.g., accuracy, precision, recall) achieved by different DL techniques in SAR ship classification?
    RQ3: What is the impact of incorporating advanced techniques such as pyramid features, feature fusion, and generative methods on the performance of SAR ship classification using DL?
    RQ4: What are the commonly employed network architectures (e.g., convolutional neural network (CNN), ResNet) for SAR ship classification using DL, and how do they compare in terms of performance?
    RQ5: What are the limitations and challenges associated with the existing DL techniques for SAR ship classification, as identified in the literature?
    RQ6: Are there any specific datasets or pre-processing techniques commonly used in the studies on SAR ship classification using DL?
These research questions aim to provide insights into the utilization, performance, challenges, and future directions of DL methods for SAR ship classification.
II-B Search Query
Search strings are crucial tools for conducting effective research. They help researchers filter through vast amounts of information to identify relevant and accurate sources. By crafting a well-defined search string, researchers can increase the precision of their search results, saving time and effort in the research process. To retrieve accurate and relevant results focused on DL techniques for SAR ship classification, the following search string was formulated:
(Deep learning OR Neural network) AND (Ship classification OR Vessel classification OR Ship categorization) AND (Synthetic Aperture Radar OR SAR)
This search string aims to capture studies that specifically address the utilization of DL or neural network approaches in SAR ship classification tasks and includes variations of terms such as deep learning, neural network, ship classification, vessel classification, ship categorization, synthetic aperture radar, and SAR.
II-C Study Selection
To identify a relevant set of papers for the systematic literature review, the study selection process involved the following steps:
Step 1: Initial Search
An initial search was conducted using the search query on a range of established academic databases and search engines, including Web of Science, IEEE Xplore, Scopus, Springer, Science Direct, ResearchGate and Google Scholar. This comprehensive approach ensured a thorough examination of the existing literature on the topic.
Step 2: Screening
Two independent reviewers screened the search results by examining the titles and abstracts of the identified papers. Inclusion and exclusion criteria were applied at this stage to determine the relevance of each paper to our research topic.
Step 3: Full-Text Assessment
The papers that passed the initial screening were retrieved in full-text format. Two independent reviewers assessed the full texts of these papers to determine their eligibility for inclusion in the systematic literature review. Any disagreements were resolved through discussion and consensus.
Step 4: Data Extraction
The included papers underwent data extraction, where relevant information was systematically extracted. This included details such as author(s), publication year, objectives, methodology, evaluation metrics, and key findings.
Step 5: Quality Assessment
The quality of the included papers was assessed using appropriate quality assessment tools or checklists. This step aimed to evaluate the methodological rigor and reliability of the selected papers.
Step 6: Data Synthesis
Finally, the data extracted from the included papers were synthesized and analyzed to derive meaningful insights, identify trends, and answer the research questions outlined in our systematic literature review.
The study selection process was conducted meticulously to ensure the inclusion of high-quality and relevant papers that contribute to the body of knowledge on SAR ship classification using DL. The inclusion criteria that were applied during the study selection process were inspired by [18], which derives them from the PICOC approach (Population, Intervention, Comparison, Outcome, Context) proposed by [19] to ensure a focused and comprehensive selection.
II-D Inclusion Criteria
This survey includes papers focused specifically on ship classification from satellite SAR images using DL techniques, algorithms or models, and appearing in peer-reviewed journals, conference proceedings or other reputable scholarly sources. Additional criteria to be satisfied were
  1. 1.
    Studies reporting performance metrics, evaluation results, or quantitative measures of DL model effectiveness.
  2. 2.
    Studies published in English to ensure accessibility and facilitate data extraction and analysis.
  3. 3.
    Studies using publicly available datasets.
  4. 4.
    Studies published after 2010.
All the papers not complying with these requirements were excluded from any further analysis.
III Results
III-A Search and screening Results
The search query was executed across multiple databases; the papers were retrieved from each database through the mentioned query string, Section II-B, with no further selection based on relevance and inclusion criteria. The numbers of retrieved titles are detailed in Table I.
TABLE I: Number of Papers Retrieved from Different Databases
Database Number of Papers
Web of Science 52
IEEE Xplore 50
Scopus 63
Springer 87
Science Direct 30
Research Gate 100
Google Scholar 1450
The initial screening reduced the total number of papers from 1832 to 161 by excluding duplications, irrelevant search results and papers that fell outside the defined inclusion criteria. Figure 1 illustrates the distribution of papers remaining after this assessment, specifying their provenance by database.
WOSIEEEScopusSDGS0020202020404040406060606080808080100100100100333333332323232319191919222284848484DatabaseNumber of PapersDistribution of Papers Across Databases
Figure 1: Distribution of titles surviving the initial screening.
During the full-text assessment stage, the content of the remaining 161 papers was examined to determine their relevance to research questions and inclusion criteria. The aim of this analysis was to identify the most pertinent studies contributing significantly to the development of DL-based SAR ship classification.
By rigorous evaluation of the full texts of these papers, it was ensured that only high-quality, relevant, and valid studies were included in the final selection. This process is essential for maintaining the integrity and reliability of this systematic literature review and for extracting meaningful insights to address the research objectives. The full-text assessment further refined the set of papers, reducing them to a final number of 49 (see Table II) and allowing for the derivation of well-founded conclusions and evidence-based findings.
TABLE II: Comprehensive List of Relevant Research Papers
Year Study
2017 [20]
2018 [21], [22], [23]
2019 [24], [25], [26], [27], [28]
2020 [29], [30]
2021 [31], [32], [33], [34], [35], [36], [37], [38], [39], [40], [41]
2022 [42], [43], [44], [45], [46], [47], [48], [49], [50], [51], [52]
2023 [53], [54], [55], [56], [57], [58], [59], [60], [61], [62], [63], [64]
2024 [65], [66], [67], [68]
After the meticulous and thorough evaluation of the full texts described above, we derived the taxonomy illustrated in Figure 2, which shows models, datasets, augmentations and learning techniques used in the literature for SAR ship classification.
Figure 2: DL Taxonomy in SAR Images for Ship classification: The ”*” in CNN denotes convolutional neural networks that were specifically designed by researchers rather than using predefined architectures.
III-B Survey Findings
The analysis of relevant literature revealed several key insights into the utilization of DL techniques for ship classification on SAR images. Here is a list thereof, ordered by research questions (see Section II):
  1. RQ1
    Evolution of Deep Learning Techniques: The adoption of CNN models initially stemmed from their success in other image processing tasks. However, as the demand for enhanced performance intensified, researchers integrated various image processing techniques and handcrafted features to augment the effectiveness of DL models [31, 68, 67].
  2. RQ2
    Performance Characteristics: Accuracy is the most commonly used performance metric in the literature for evaluation. However, its widespread use highlights the need for incorporating additional metrics, particularly when dealing with long-tailed, imbalanced datasets. Relying solely on accuracy can be misleading in such scenarios, as it may not adequately capture the model’s performance across different classes, especially for minority classes.
  3. RQ3
    Impact of Advanced Techniques: Techniques like pyramid features, feature fusion, and transfer learning were employed to bolster ship classification performance on SAR images [40, 64, 58, 59, 46, 25]. Handcrafted features and pyramid features demonstrated particularly promising results. However, the research in this area remains in its nascent stages, warranting further exploration of these advanced techniques.
  4. RQ4
    Commonly Employed Network Architectures: Specific CNN architectures, particularly deeper models like VGG and ResNet, have shown promise in SAR ship classification due to their ability to extract complex features from SAR imagery. Interestingly, simpler and shallow CNN models, when combined with effective image processing techniques like noise reduction and image enhancement, have also achieved commendable accuracy.
  5. RQ5
    Limitations and Challenges: SAR ship classification faces several critical challenges. The scarcity and imbalance of labeled SAR image data, coupled with limitations in SAR image processing techniques, hinder model performance. Additionally, relying solely on DL networks designed for optical imagery, or choosing inappropriate SAR-specific architectures, can lead to suboptimal results. Furthermore, SAR-AIS matching, while valuable, has limitations due to its inherent unreliability (e.g. lack of transmission, noisy and inconsistent messages [69]). The development of robust, automated SAR-AIS matching algorithms is crucial. Finally, SAR processing models can benefit significantly from tailored loss functions [23, 34, 48] that effectively filter out irrelevant information and focus on the features of target objects. Research on these tailored loss functions is an active area of development.
  6. RQ6
    Specific Datasets and Pre-processing Techniques: Studies in the literature commonly utilize two publicly available datasets: FUSAR-Ship [17] and OpenSARShip [70]. Geometric pre-processing techniques, such as cropping, rotation, flipping, and scaling, are frequently employed to improve the generalizability of ship classification models. However, the necessity for these techniques arises from the challenges associated with small and low-resolution vessel representations in the data. Super-resolution techniques offer a potential avenue for addressing these limitations [71, 72, 73]
IV Background
Several techniques are employed for SAR ship classification using DL, as illustrated in Figure 2. This section is divided into four parts. The first part explores the various DL techniques utilized in the literature for SAR ship classification. The second part examines the impact of handcrafted features on the performance of DL models. As SAR images possess unique attributes, the third part discusses how these attributes are leveraged in SAR ship classification. Finally, the fourth part investigates the use of fine-tuning for SAR ship classification.
IV-A Deep Learning Techniques
As apparent from Table II, we found the origin of SAR-based ship classification using DL in 2017, when [20] classified 4 different ship types among the targets detected in 6 VV-polarized COSMO-SkyMed scenes using a simple CNN model with 2 convolution layers and 2 fully connected layers; the images were cropped to $64\times 6464 × 64$, 4 scenes were used for training, 1 for testing and 1 for validation. Their results showed that the model was able to learn the cargo ships better than the others due to their high prevalence in the training set. The advancement of SAR platforms over time has resulted in making images with better and better spatial resolutions available. [21] treats images from TerraSAR-X, with resolution of 3 meters per pixel, using CNNs. They prove that by using multiple-resolution SAR images, CNN models are good at classification tasks, achieving an f1-score of 94%. Four augmentation techniques were applied to form the training set: reflections (horizontal and vertical), rotations (-10 to 10 degrees), translation (-5 to 5 pixels) and noise addition (randomly changing pixel values to zero).
[23] explores two important aspects of a DL model: attention based CNN model and a novel loss function for class-imbalanced datasets named “weighted distance measure”. A VH-polarized data set with Ground Range Detected (GRD) images is divided into 4 classes, i.e. cargo, tanker, tug and others. The results show that the model is capable of learning class representation, achieving a recognition accuracy of 84%.
[27] evaluates a convolution network for detection and classification. The input data are drawn from the PALSAR-2 L-band SAR with 10 m resolution in HV polarization, and the results indicate that the CNN-based models can detect ships, but the classification performance is affected by ships that look alike.
Ship classification can also be achieved using a sequence-based approach. [29] constructs a sequence input by capturing ten consecutive images of the same target at different times, degree of rotations. The authors employ a combination of CNN and long short-term memory (LSTM) with 4 categories of OpenSARShip data (boat, cargo, container, and oil tanker). Their findings suggest that the model learns improved class representations when utilizing image sequences.
A CNN, when not fully connected, can be used as a feature extractor. The extracted features can then be fed to a classifier or further processed using methods like metric learning. Metric learning involves training a model to measure the similarity or distance between data points, often to cluster extracted features in a way that makes them distinguishable, thereby enhancing the model’s performance. [30] compares the performance of a CNN-only approach to a CNN combined with metric learning, on OpenSARship (bulk carrier, container, and oil tanker) focusing on data with larger sample sizes, then enhances the features extracted by the CNN by concatenating additional features learned using metric learning. The results show that the CNN combined with metric learning outperforms the simple CNN.
[57] proposes a novel Hierarchical Embedding Network with Center Calibration (HENC) for few-shot ship classification. HENC utilizes a CNN to extract hierarchical features at different scales. This method achieves an accuracy of 81.03% on the six-category (cargo, tanker, tug, fishing, dredging and others) OpenSARShip dataset with few-shot training, demonstrating the effectiveness of deep models in learning class representations even with limited data.
[58] discusses a deep metric learning approach that leverages a pairwise representation fusion for multi-proxy learning, and introduces a novel distribution loss to refine the sample distribution. This allows for both clustering similar classes together and training the model to make a clear distinction between all classes. The effectiveness of this three-pronged approach is evaluated on two optical ship datasets (FGSC-23 and FGSCR-42) and one SAR dataset (5-class FUSARShip).
[64] introduces HDSS-Net, a novel deep learning architecture for ship classification in SAR images. This addresses challenges such as large variations in ship size, the elongated shapes of ships in SAR images, and intraclass diversity alongside interclass similarity. HDSS-Net employs a hierarchical design with three key modules. The first, a feature aggregation module (FAM), creates a multi-scale feature pyramid. The second, a Feature Boost module (FBM), utilizes rectangular convolutions to capture the elongated features of ships. Finally, a spherical space classifier (SSC) processes features in a spherical space to enhance class separability. Evaluations on OpenSARShip (three and six classes) and FUSAR-Ship (seven classes) datasets demonstrate that HDSS-Net outperforms traditional CNN methods, highlighting its potential value for maritime surveillance applications.
There exists a trade-off between spatial resolution and the cost of SAR imagery. Developing algorithms that achieve accurate results using medium-resolution data is essential for broader applications. [34] uses DenseNet and metric learning to perform SAR ship classification on three-class (tanker, container, bulk carrier) and five-class (tanker, container, bulk carrier, cargo, general cargo) datasets from OpenSARShip and proposes a novel loss function to incorporate deep metric learning. To help the model learn accurate class representations, it utilizes images in triplets [74]. Furthermore, data augmentation techniques such as flipping, rotation, translation and Gaussian noise are applied to address overfitting. The results demonstrate superior performance in both the three-class and five-class categories.
Each CNN layer stores important information about features, but some information may be lost when features pass from one layer to another. [33] proposes a novel multi-scale CNN (MS-CNN) method to utilize all the information from each convolutional layer. Using three classes from the OpenSARship dataset (bulk carrier, container ship and tanker), the MS-CNN, integrated with information from each convolutional layer, achieves a 5% better performance than the same CNN without integration. This improvement suggests that information loss occurs within the network.
[54] uses a deep network with a novel inception-residual controller (IRC) module to improve the performance. The authors train the model on a five-category dataset from FUSARShip. The data underwent homogenization, normalization, and augmentation techniques such as flipping and rotation and, after 50 training epochs, the model with IRC achieves an accuracy of 98.71%, indicating that deeper models trained on pre-processed data can significantly improve the classification performance.
[36] proposes a novel multi-scale dual-attention CNN that leverages information from convolutional layers. The authors employ a VGG-Net architecture for feature extraction and utilize a three-class dataset (cargo, container ship, and tanker) from OpenSARShip. The methodology involves modifying the VGG network to extract features from SAR images, followed by the application of squeeze-and-excitation (SE) blocks to focus on both channel-wise and positional context. SE blocks learn the importance of feature maps, promoting informative features and suppressing less useful channels. The results demonstrate the effectiveness of this method, achieving an accuracy of 92.66%. This success suggests the potential of attention mechanisms in this field. [59] proposes an enhanced vision transformer model that incorporates feature fusion modules to capture local information, followed by SE attention. The effectiveness of this approach is evaluated on a three-class OpenSARship dataset, achieving an accuracy of 76.25%.
The convergence in training a DL model is controlled by several parameters, among which the learning rate is of paramount relevance. It determines how the model adjusts its internal parameters during training, and can be either fixed throughout the process or adjusted dynamically. [47] explores three learning rate management strategies and their impact on transfer learning. Findings suggest that a smooth learning rate change can improve the model’s ability to understand the data, leading to enhanced classification performance. Additionally, the authors investigate the effect of freezing different layers during transfer learning on a CNN, emphasizing the importance of layer analysis before fine-tuning. The models are first trained on the ten-category MSTAR dataset (SAR images of tanks, [75]) and then fine-tuned on a three-category OpenSARShip dataset (cargo, container ship, and bulk carrier). The results demonstrate that incorporating learning rate scheduling and thoughtful layer freezing significantly improves the model performance. This finding underscores the importance of employing effective training strategies for optimal SAR ship classification.
Selecting optimal DL architectures for specific tasks can be challenging. Traditionally, researchers rely on exploring existing architectures and their applications. However, this approach often lacks a systematic methodology, leading to difficulties in finding the ideal architecture for a given classification problem. To address this challenge, [50] proposes a lightweight searched binary neural network (SBNN), obtained through network architecture search (NAS), quantized for weight reduction and incorporating patch shift processing to remove noise. The authors employ three (bulk carrier, container and tanker) and six-class (bulk carrier, container ship, tanker, cargo, general cargo and fishing) datasets from OpenSARShip, containing both VV and VH polarizations. They apply data augmentation methods such as random cropping and horizontal flipping. The SBNN achieves accuracy scores of 80.03% and 56.73% for three and six classes, respectively. NAS not only results in superior performance but also lead to a lightweight model. [53] introduces a novel mini-size searched convolutional metaformer (SCM) that leverages an updated NAS algorithm with progressive data augmentation (crop, flip, rotate, resize, cutout) to discover an efficient convolutional network baseline. This baseline is then enhanced with a transformer classifier to improve spatial awareness, and further augmented with a ConvFormer cell, combining the searched convolutional cell within a Metaformer block for superior feature extraction. Evaluated on VV- and VH-polarized datasets from OpenSARShip (3 categories, i.e. bulk carrier, container, and tanker) and FUSARShip (4 categories, i.e. bulk carrier, fishing, cargo, and tanker), SCM achieves accuracies of 82.06% and 63.9%, respectively, with a lightweight design (about half a million weights). [48] proposes a one-shot learning approach using a siamese network. The approach employs a 7-layer fully connected feature fusion that enables the network to classify 16 categories within the OpenSARShip dataset. This methodology involves three stages: (i) preprocessing the images and converting them to three dimensions by adding two additional channels (actual, filtered, and inverted SAR Image), (ii) creating image triplets containing an anchor image, a positive image, and a negative image, and (iii) training the model using a loss function that ensures the anchor image to be closer to the positive image and farther from the negative image. The results demonstrate the model’s effectiveness for ship classification.
[61] addresses the challenge of handling the long-tailed distribution problem in imbalanced SAR data. The authors propose a novel architecture that incorporates a multi-branch expert network with a dual-environment sampling strategy. This network consists of three classifiers trained to handle different distribution scenarios: long-tailed, balanced, and inverse long-tailed. Intraclass dual-environment sampling adjusts sampling rates within each class based on classification confidence, creating varied distributions. Combined with center loss [76], this method enhances robustness and accuracy, outperforming existing techniques on the MSTAR and FUSAR datasets. Experiments confirmed significant improvements in recognizing both majority and minority classes, demonstrating the effectiveness of the proposed approach in balancing feature learning and improving overall and average accuracy.
Due to the resolution of SAR images, research on small-sized fishing vessels has not been given sufficient importance. However, as the resolution of SAR systems increases, there is a growing need to address the classification of fishing vessels. [62] proposes a novel architecture for classifying fishing vessels, consisting of three modules: (i) the multi-feature extraction (MUL) module, which extracts features using DenseBlocks and channel-wise concatenation, (ii) the feature fusion (FF) module, which aggregates features from different layers, and (iii) the parallel channel and spatial attention (PCSA) module, which applies channel and spatial attention. The paper also proposes a method for SAR and AIS integration to create a fishing vessel dataset. This methodology was tested on three fishing vessel classes, and the model achieved an accuracy of 89.7%.
IV-B Handcrafted Features
Traditional features like histogram of oriented gradients (HOG) can be combined with features extracted by a CNN to enhance the performance of DL models that require a large amount of features. HOG features capture the distribution of edge orientations within an image region, describing its local shape and appearance. [31] uses a VGG network as the backbone for feature extraction and concatenate these features with HOG features on three categories (bulk carrier, container ship and tanker) from the OpenSARship dataset. The results indicate that the model that incorporates HOG features achieves 7% better performance than the model without HOG features.
[43] proposes a novel method called HOG-ShipCLSNet. This method first extracts HOG features and then learns deep features using a CNN called the multiscale classification mechanism (MS-CLS). MS-CLS extracts features from each convolutional layer. Since features from different layers have varying dimensions, the network employs a global self-attention mechanism to focus on important features within each layer. To ensure stable training, the fully connected balance mechanism resizes features to a uniform size and then averages them. Finally, the HOG feature fusion mechanism merges the handcrafted HOG features with the DL features before performing classification. The authors conducted experiments on two datasets: three categories from OpenSARShip (bulk carrier, container ship, tanker) and seven categories from FUSARShip (bulk carrier, container ship, fishing, tanker, general cargo, other cargo, other). The results indicate that HOG-ShipCLSNet is effective for ship classification. [67] proposes a novel attention-guided neural network for ship classification using the OpenSARship dataset. This network incorporates 14 handcrafted features processed by a feed-forward network and combines them with features extracted by the attention-guided network. These combined features are then employed for training a classification model. The model, trained on 4 categories from OpenSARship, is evaluated on a dataset of SAR-AIS integrated Sentinel-1 SAR images. This approach achieves an accuracy of 86.3%, demonstrating the potential effectiveness of combining handcrafted features with an attention mechanism for ship classification tasks.
Choosing effective features plays a vital role in classification performance: meaningful features help the model to be trained optimally. [60] compares four different handcrafted features (Canny, Harris, Gabor, and local binary patterns (LBP)) and proposes a novel method for feature fusion that optimizes this process. The method consists of two blocks: (1) the knowledge supervision and collaboration module (KSCM), which facilitates mutual learning between handcrafted and deep features, and (2) the feature fusion and contribution assignment module (FFCS), which enhances channel attention. The proposed model has been tested on three categories from OpenSARShip and seven categories from FUSARShip using both deep and shallow versions of AlexNet, VGGNet, ResNet, and DenseNet. The results demonstrate that handcrafted features enhance the performance of all models, with Gabor features showing the most significant average feature gain of 3.4%. Additionally, shallow networks (VGG-11 and ResNet-19) exhibit the largest improvement over their baseline counterparts following feature fusion.
[37] investigates the effect of combining handcrafted features with DL features by considering four handcrafted features (HOG, naive geometric features (NGF), local radar cross section (LRCS), and principal axis features (PAFs)) in conjunction with four DL models (AlexNet, VGGNet, ResNet, and DenseNet). They analyze different feature injection locations and techniques, recommending injection in the final fully-connected layer with dimension unification weighted concatenation and feature normalization (DUW-Cat-FN) for optimal results. The authors utilize three classes (bulk carrier, container ship, and tanker) from OpenSARShip and seven classes (bulk carrier, container ship, fishing, tanker, general cargo, other cargo, and other) from FUSARShip. The results reveal that injecting handcrafted features into a DL model has the potential to improve performance by up to 6%, thus highlighting the continued importance of handcrafted features. [68] proposes a novel contrastive learning-based method that leverages the fusion of handcrafted features with SAR images. It employs data augmentation techniques on both SAR images, and handcrafted features, applying distinct augmentations to each data type. Subsequently, the model calculates three types of losses during deep training to facilitate learning of two primary pretext tasks and a cluster-level task. The proposed model, named DCPNet, achieves accuracies of 73.66% and 87.94% for 3-class and 7-class classifications on the OpenSARship and FUSAR-Ship datasets, respectively. These results suggest the potential effectiveness of contrastive learning for ship classification tasks.
Images in OpenSARship contain information such as the length and breadth of the vessel. [32] enhances the ship classification performance of a DL model by combining the output of a CNN and a KNN (K-nearest neighbour classifies new data based on the majority class of its nearest neighbors). The KNN is trained to classify ships based on length and breadth taken from AIS data, while the CNN uses the image data. Three classes (cargo, tanker, and others) from OpenSARship dataset are used. The model combined with KNN showed superior performance compared to the CNN model alone.
Since a ship comprises various sections, [41] first employees a minimum bounding rectangle (MBR) to detect the ship in the image. Then, subdivide the bounding box into the bow, midship, and stern sections and extract the features using two models based on the VGG network. One model leverages the data from the three sections to capture local features, while the other analyzes the entire MBR to capture global features. Finally, feature fusion is employed to classify the ships into two categories: military and civilian. The method, applied to high-resolution images captured by GaoFen-3, achieves an accuracy of 96%.
IV-C Use of SAR properties
Maritime monitoring represents a unique challenge in case of moving ships. Indeed, motion can cause the ship to appear shifted within the SAR image with respect to its actual geographical position. This information can be proficiently exploited for classification purposes. [35] proposes a complex-valued CNN (CV-MotionNet) that considers both amplitude and phase information from single-look complex SAR images. The experiments were conducted on two datasets: a five-class simulated dataset and a three-class real-world GaoFen-3 dataset containing cargo ships, tankers, and other categories. The results demonstrate the model’s effectiveness in classifying SAR targets.
Sidelobes, faint echoes in SAR images caused by the finite length of the radar pulse, can hinder the performance of DL models for ship classification. [38] investigates the effect of sidelobe elimination on ten classes (anti-pollution, bulk carrier, container, dredger, fishing, cargo, law enforce, passenger, oil tanker, and port tender) from FUSARShip using a simple CNN model. The results show that supressing the sidelobes improves the model performance, achieving up to 98% accuracy, which indicates the effectiveness of sidelobe elimination.
[40] investigates whether using information from two different polarizations (VV and VH) can improve the model performance by capturing more features. The authors propose a novel network called squeeze-and-excitation Laplacian pyramid network with dual-polarization feature fusion (SE-LPN-DPFF). This network first merges features extracted from both polarizations. Then, it utilizes squeeze and excitation blocks to enhance feature representation by selectively emphasizing informative features. Finally, the features are fed into a Laplacian pyramid network based on residual blocks, which extracts information from different resolutions and performs classification ensuring that the model is trained on this fine-grained information. The authors employ two datasets from OpenSARship: one with three categories (bulk carriers, container ships, and tankers) and another with six categories (bulk carriers, cargos, container ships, fishing, general cargos and tankers). To ensure a balanced dataset, they use 70% of the data from the category with the fewest samples as the training set and the remaining data for testing. They compare the performance of SE-LPN-DPFF to 35 other DL classifiers. The results show that SE-LPN-DPFF outperforms all other models.
[46] proposes a different approach using a dual-polarized network based on a pretrained VGG architecture. This network learns representations from both VV and VH polarizations and then apply bilinear pooling. They train the model on a three-class dataset (tanker, container ship, and bulk carrier) from OpenSARShip, where VV and VH images are paired for training. The results show that the model outperforms state-of-the-art models, achieving an accuracy of 88%.[49, 56] propose a novel methodology combining a loss function and bilinear pooling, termed the group bilinear convolutional neural network (GBCNN), based on dense blocks. Grouped VV and VH images are used to train two parallel models. The extracted features are then passed through bilinear pooling. The resulting features from both pooling layers are sent to the loss function, along with a third feature created by merging the pooling outputs. Utilizing three- and five-class (tanker, container ship, bulk carrier, cargo, general cargo) datasets from OpenSARShip with data augmentation techniques such as flipping, rotation, translation, and Gaussian noise, the GBCNN achieved accuracy scores of 88% and 67% for three and five classes, respectively.
[51] investigates the polarization of images in conjunction with handcrafted features for ship classification. It proposes a novel polarization fusion-based geometric features network (PFGFE) that utilizes geometric features alongside VH-polarized, VV-polarized, and a combined VV-VH image. PFGFE employs three VGG network-based CNNs for feature extraction. A global feature self-attention mechanism is then applied to the output of each CNN. Subsequently, the extracted features are merged, and their classification probabilities are concatenated with corresponding coefficients. Geometric features are processed using LSTM to capture the importance of long-term dependencies. Finally, the outputs from the CNNs and LSTM are concatenated, and this final representation is used for classification. The authors evaluate PFGFE on three-class (bulk carrier, container ship, and tanker) and six-class (bulk carrier, cargo, container ship, fishing, general cargo, and tanker) datasets from OpenSARShip. The results demonstrate that PFGFE achieves accuracies of 79.84% and 56.83% respectively.
Building on the idea that exploiting dual polarization information can enhance performance, [55] proposes a novel dual-polarization information-guided network (DPIG-Net) that leverages a polarization channel cross-attention framework (PCCAF) and a dilated residual dense learning framework (DRDLF) for effective feature extraction and fusion. It evaluates DPIG-Net on the OpenSARShip dataset using VH and VV polarizations for three (bulk carrier, container ship, tanker) and six classes (adding cargo, fishing, and general cargo), achieving accuracies of 81.28% and 58.68%, respectively. Smoothed polarization fusion, coupled with attention modules such as multiscale feature fusion and dynamic gating feature fusion, has been shown to improve model accuracy for ship classification. [77] employs this approach on OpenSARShip, achieving accuracies of 87.13% and 65.97% for three and six classes, respectively. These results suggest that smoothed polarization fusion may be more effective than coherence fusion [55]. [66] proposes a novel dual branch deep network for ship classification, tested on OpenSARship. The network leverages VH and VV polarizations and comprises three modules: (1) an image construction module that generates two pseudo-RGB images representing polarimetric and texture features, (2) a feature extraction module that employs two pre-trained ConvNeXt networks to extract features from the constructed images independently and (3) a feature fusion and classification module that combines the extracted features and performs classification. This network achieves accuracies of 87.62% and 64.26% for 3-class and 6-class classification tasks, respectively.
The limited spatial resolution of SAR imagery presents a significant challenge for ship classification tasks, particularly for small vessels. These ships often occupy only a few pixels within the image, making it difficult to extract the features needed for an accurate classification. To address this, [26] proposes a joint convolutional network for enhancing the resolution of SAR images before classification. It uses TerraSAR-X images for pre-training and three classes (bulk carrier, container, tanker) from OpenSARship for training. The data are first fed into a generative model for upscaling, and then classified using a dense blocks-based convolutional network. It also compares the performance of models trained on VV and VH polarizations. The results show an increase in performance compared to simpler networks, achieving classification accuracies of 81.02% and 79.12% for VH and VV, respectively.
When data are scarce and imbalanced, some DL methods, such as the siamese networks, can be helpful in overcoming this challenge. Siamese networks focus on learning a similarity function by training two networks that share weights, allowing for better learning of data representations. [28] proposes a novel loss function and apply a siamese network architecture. It utilizes two inception blocks on different polarizations (VV and VH) for four classes (cargo, tanker, tug, and other) from OpenSARship. The images are first passed through a target recognition algorithm (RCNN) to crop the targets. Then, the cropped images are fed into the inception blocks, the features are extracted and the loss is calculated. The results indicate that siamese networks with target detection outperform models without target detection, achieving an accuracy of 87.04%.
In ship classification tasks, researchers typically use datasets with three or four categories. This is because DL models require a large amount of data for training, and maritime data are inherently imbalanced, making it challenging to achieve good performance with more classes. However, a study in [42] demonstrates the potential of feature fusion for handling a larger number of categories. It employees an eight-category dataset (cargo, tanker, carrier, container, fishing, dredger, tug, and passenger) from OpenSARShip. A VGG-based feature extractor is used to simultaneously extract feature maps from both VV and VH polarizations. These feature maps are then merged. Finally, a novel hybrid loss function was applied to reduce intra-class distances and increase inter-class distances, leading to improved classification performance. The results, with a mean average precision of 82%, support the effectiveness of this approach.
IV-D Fine-Tuning
High resolution images from the CosmoSkyMed mission (level 1A HIMAGE single-look complex Slant (SCS) product images in the X-band with single HH or VV polarization) are exploited to test the effectiveness of pre-trained models on a small dataset with three classes. [22] conclude that fine-tuning a VGG16 network can lead to f1-scores of up to 97.8%. This shows the usefulness of DL techniques for datasets with less samples like SAR.
[24] demonstrates that leveraging transfer learning with pre-trained models, combined with effective data augmentation strategies, can significantly improve the classification accuracy even when working with small training datasets, achieving an accuracy of 98% using a deep architecture, DenseNet-121. The input images are sourced from the TerraSAR-X stripmap mode and are manually annotated by interpretation experts using AIS data. There are only three classes: bulk carrier, container ship, and oil tanker. The data augmentation techniques applied are flipping, brightening, sharpening, random cropping, and rotating.
SAR data acquisition involves varying incident angles. [25] examines the effect of these varying angles on model performance. It uses a dataset containing images of three classes taken from Cosmo-SkyMed and GaoFen-3, encompassing incident angles ranging from 20 to 60 degrees. The data were divided into three overlapping subsets: one containing the entire dataset, another containing data with low incident angles (less than 48 degrees), and a third one containing data with high incident angles (greater than 45 degrees) . All three datasets were tested on four pre-trained DL models: VGG-16, VGG-19, ResNet-50, and ResNet-101. The results indicate that incorporating data from all available angles enhances the models’ predictive accuracy.
Data scarcity can be addressed by training the model on domain-specific data, such as optical images of ships, and then applying domain adaptation to SAR data. [44] proposes a novel method, DSAN++, for domain adaptation. This method involves first training on optical images and then transferring the learned knowledge to SAR data. DSAN++ utilizes two ResNet networks for feature extraction. The extracted features undergo feature fusion, followed by the application of a subdomain adaptation network to learn transferable representations. To capture both local and global distributions, shallow and deep ResNet models (ResNet-18 and ResNet-50) are integrated with the convolutional block attention module. The experiments were conducted by training the model on a four-category (cargo, bulk carrier, container ship, and oil tanker) optical remote sensing (ORS) dataset obtained from Google Earth. Subsequently, domain adaptation was performed on two datasets: (1) GF-SAR with three classes (bulk carrier, container ship, and oil tanker) collected from FUSAR and GaoFen, and (2) a three-class dataset (cargo, container ship, and oil tanker) derived from TerraSAR-X. The results demonstrate that DSAN++ leads to improved performance in subdomain adaptation. [52] compared the fine-tuning performance of models pre-trained on domain-specific versus non-domain-specific datasets using CNN’s. The methodology was evaluated on three-class targets (carrier, container, and oil tanker) from VV polarized TerraSAR-X, and four-class targets (cargo, bulk carrier, container, and oil tanker) from FUSARShip. The results suggest that shallow models with feature refinement can achieve better performance when fine-tuned on domain-specific models. The authors investigate the pre-training effects of ImageNet and an ORS dataset with data augmentation techniques (scaling, flipping, and rotating). While ImageNet, due to its large size, yields better performances on deep networks, the ORS dataset enhanced the performance of shallow networks, indicating that shallow models are more effective with less data.
IV-D1 Data Augmentation & GANs
Techniques like data augmentation, when combined with transfer learning, can address data-related problems. [39] demonstrates that using transfer learning with data augmentation and manual features can improve the performance of a DL model. It uses ground range detected SAR data acquired in interferometric wide-swath mode from OpenSARShip. The data consist of two channels, with VH and VV polarizations, stacked together. The study focuses on classifying ships into four categories: bulk carrier, container ship, fishing vessel, and tanker. The authors use five data augmentation techniques (horizontal flipping, vertical flipping, rotation by 90°, rotation by 270°, and Gaussian noise injection) and incorporate 14 hand-crafted features. The effectiveness of these hand-crafted features combined with DL features are compared across 3 machine learning (Radial Basis Function (RBF)-SVM with two stage feature selection, SVM classifier combination, SVM one-vs-rest) and 4 DL models (CNN, AlexNet, VGG19, and ResNet50). Additionally, the authors present a methodology for matching AIS and SAR data and test the performance on real data collected from three different locations. Their results show that the pre-trained ResNet50 model, when fine-tuned on SAR images and hand-crafted features, achieves an accuracy of 82.5%.
One of the possible DL-based solutions to the data scarcity problem in optical images is represented by generative models and can also be applied in SAR images. These models learn data representations and produce images that resemble the originals, effectively increasing the available data. [45] proposes a novel conditional generative adversarial network (GAN) method that generates high-quality SAR images, improving ship classification performance. The method consists of two parts: (1) a residual convolutional block is integrated with the generator to enhance the detail of generated images, and (2) the generated images are merged with real images and sent to a discriminator for two tasks: distinguishing real from fake images and classifying them into three classes. The authors use a three-class dataset (cargo, container, tanker) from high-resolution SAR images of the Chinese C-band, 3-meter resolution GaoFen-3 platform. The results indicate that the GAN is capable of generating high-resolution images, and that these images, coupled with the real images, enhance the classification performance.
[65] proposes a generative model-based domain adaptation framework for ship classification on a 5-class FUSARShip dataset. This framework leverages domain adaptation by transferring knowledge from the optical images of the FGSCR42 dataset. [63] introduces the Attention-Dense CycleGAN (ADCG) to improve recognition by leveraging transfer learning from optical to SAR domains (OPT2SAR). ADCG enhances CycleGAN with a dense connection module (DCM) and a convolutional block attention module (CBAM). Using FUSARShip and the FGSCR optical dataset, ADCG achieves superior performance, evaluated via Fréchet inception distance (FID) and kernel inception distance (KID). It demonstrates a 6% improvement in recognition accuracy for SAR ship classification networks, thus validating its effectiveness in addressing class imbalance and sample scarcity through effective transfer learning.
V Findings
V-A Datasets
The literature explores two public datasets, FUSARShip and OpenSARShip, and some additional, distinct data sources for SAR ship classification. Their detailed information is provided below.
V-A1 OpenSARShip
The OpenSARship dataset is a publicly available resource designed to support research in ship classification using SAR imagery. This dataset provides a collection of pre-processed SAR image chips specifically focused on ship detection and classification tasks. The current version contains 34528 vessel instances. The data are collected using two products from ESA’s Sentinel-1 repository (single look complex, and ground range detected) in interferometric wide swath mode. The dataset consists of 16 ship classes, containing classes like Cargo, Tanker, Fishing, Tug, Dredging. However, the dataset suffers from class imbalance, with 61% data belonging to Cargo category. Three classes represent 95% of data, while the remaining 13 classes only represent a 5% of the data provided. This imbalance can lead any model to bias toward the majority class. The literature shows that researchers often address this problem by only using the 3 to 5 most populated classes, e.g., cargo, tanker, tug, fishing, and dredging [23, 29, 57, 34, 36, 43, 60], or by utilizing subclasses within the cargo category, e.g., general cargo, container ship, bulk carrier [30, 33, 47]. OpenSARship data incorporates polarization information, VH and VV, which researchers can leverage to create fusion networks [50, 53, 40, 46, 49, 56, 51, 55, 66]. The dataset’s main weaknesses such as low object resolution (20 m/pixel) and class imbalance, limit the performance of a DL model. Nevertheless, if these limitations are addressed, OpenSARship serves as a valuable starting point for researchers to explore their methodologies before acquiring commercial SAR imagery.
V-A2 FUSARShip
The FUSARShip dataset provides a valuable resource for researchers in SAR ship classification. This dataset leverages high-resolution SAR images captured by China’s GaoFen-3 (GF-3) satellite, specifically designed for ocean remote sensing and marine monitoring tasks. FUSARShip is constructed from a collection of image chips extracted from over 126 GF-3 scenes, encompassing 15 primary ship categories, e.g. cargo, tanker, fishing, and 98 subcategories. Each ship instance has a size of $512\times 512512 × 512$ pixels, with a pixel size of 0.5 meters. Similar to OpenSARship, FUSARShip presents a challenge due to class imbalance. Three subcategories constitute approximately an 81% of the data, while the remaining 95 categories collectively only represent a 19% thereof. This significant imbalance can lead to models exhibiting bias towards the overrepresented classes. Consequently, researchers often address this issue by focusing on the 3-5 most populated classes, e.g., cargo, fishing, tanker, bulk carrier [58, 54, 53, 44, 52, 65, 63] as observed in the OpenSARship dataset.
V-A3 Other Data Sources
Large datasets are crucial for training a DL model. However, the imbalance data limitation in the above datasets can hinder model performance. Researchers address this challenge using two techniques. Firstly, they leverage high-quality data from various satellite missions, see Table III, such as TerraSAR-X [21, 26, 24, 44, 52], CosmoSkyMed [22], PALSAR-2 [27], and GaoFen-3 [41, 35, 25, 45]. These missions not only provide high quality data, but also allow researchers to exploit AIS data to strategically select images based on specific locations and timeframes. For instance, if a researcher aims to develop a model for detecting fishing activity during a particular season in a specific region, AIS data can be analyzed to identify relevant satellite images for purchase.
Secondly, due to the cost associated with acquiring high-resolution commercial imagery, researchers often employ transfer learning techniques. This involves training a model on a large, general-purpose dataset (e.g., ImageNet), or domain specific dataset, e.g., MSTAR [75] (SAR images of tanks), optical remote sensing images of ships taken from Google Earth [78] and then on a smaller, task-specific dataset (e.g., FUSARShip). Pre-training on a comprehensive dataset equips the model with the ability to extract generic features from images. Subsequent fine-tuning on the target dataset refines these features for the specific classification task.
TABLE III: SAR Platforms and their resolutions
Platform Band Image Modes Resolution(s)
Sentinel-1 C
StripMap,
Interferometric Wide Swath (IW),
Extra Wide Swath (EW)
- StripMap: $5\times 55 × 5$ m
- IW: $5\times 255 × 25$ m
- EW: $20\times 4020 × 40$ m
Cosmo-SkyMed X
Spotlight,
StripMap,
Wide
- Spotlight: Up to 1 m
- StripMap: Up to 3 m
- Wide: Up to 30 m
TerraSAR-X X
StripMap,
Spotlight,
High Resolution Spotlight (HRS)
- StripMap: Up to 3 m
- Spotlight: Up to 0.3 m
- HRS: Up to 0.3 m
Gaofen-3 X
StripMap
Spotlight
- StripMap: Up to 2.5 m
- Spotlight: Up to 0.9 m
PALSAR-2 L
StripMap
Spotlight
- StripMap: $3\times 33 × 3$ m
- Spotlight: $1\times 31 × 3$ m
V-B Data Augmentation
Due to the limited availability of SAR images and the inherent data requirements of DL models, data augmentation techniques have been proposed in the literature to address this challenge. These techniques can improve the performance of DL models for SAR ship classification.
V-B1 Rotation
In the literature, rotation is widely employed to enhance the robustness of DL models for SAR ship classification [39, 49, 56, 54, 34, 21, 53]. It involves artificially rotating the training images by various angles to simulate real-world scenarios where objects might be oriented differently relative to the SAR sensor. A common approach is to rotate images within a range of +4545+45+ 45 to 4545-45- 45 degrees, as this range captures a significant portion of the potential ship orientations. This augmentation strategy helps the model learn features that are invariant to rotation, improving its generalization ability and reducing the risk of overfitting to the specific orientations present in the original training data.
V-B2 Flipping
The literature commonly employs horizontal and vertical flipping data augmentation techniques [50, 52, 56, 24, 39, 53, 54]. These techniques generate new images by mirroring the ships along the horizontal or vertical axes. This process simulates scenarios where ships are viewed from opposite directions, effectively augmenting the dataset and potentially improving the model’s ability to generalize to unseen orientations.
V-B3 Scaling
DL networks typically require a fixed-size input. Image scaling, a data augmentation technique that resizes images to a standard size, addresses this requirement. Scaling can be performed by increasing or decreasing image size, with larger images demanding significantly more computational resources for processing [52, 53]. In SAR ship classification, a common practice is to scale images to a size of 128x128 pixels, balancing accuracy and computational efficiency.
V-B4 Translation
In image processing, translation refers to the spatial transformation of an image by a specific offset in both horizontal and vertical directions. This essentially moves the entire image content without any scaling or rotation. In SAR ship classification, translation augmentation can be employed to artificially introduce variations in the training data [49, 21, 34, 56]. By exposing the DL model to images with slightly shifted ship positions, translation augmentation can potentially improve the model’s generalization ability. This is because the model learns to recognize ships irrespective of their small location variations within the image, making it more robust to real-world scenarios.
V-B5 Gaussian Filtering
SAR images of marine scenes are often corrupted by a strong noise component. To address this challenge, Gaussian filtering is commonly employed [34, 49, 56, 39]. This technique works by averaging the intensity values of pixels within a localized neighborhood around each pixel. This process reduces the random variations in pixel intensity caused by noise, enhancing the distinction of ship features from the background noise.
V-B6 Cropping
Naturally, SAR images contain a lot of useless information like background, and the object representations are very small, leading the model to learn non-informative details. Cropping is the process of extracting a specific region of interest containing ships and reducing the useless information like background sea. Cropping is commonly employed in literature to augment data [24, 28, 53, 20].
V-C Handcrafted Features
Whereas augmentation is important to help models learn from more data, it can also lead to overfitting. Moreover, augmenting low-quality images often produces additional low-quality data, which can confuse the model and hinder effective learning. In such scenarios, handcrafted features can provide valuable support by being fused with deep learning features within the model (see Fig. 5).
DL models offer significant advantages by virtue of their inherent feature extraction capabilities, but their performance in SAR ship classification can be hindered by data scarcity and imbalance. To address this challenge, researchers have leveraged handcrafted features. Here are some examples of features identified through our literature search:
V-C1 Histogram of oriented gradients
Histogram of Gradients (HOG, see Fig. 3) features are a powerful tool for extracting informative features from images, particularly those containing well-defined edges and shapes [31, 43, 37]. These features describe the local structure of an image by capturing the orientation and intensity of gradients.
Figure 3: HOG features of a SAR image
V-C2 Local Radar Cross-Section
Local Radar Cross-Section (RCS) features refer to characteristics extracted from an image that are related to the radar scattering properties of specific image regions. These features are also combined with additional DL features in the literature to improve the performance of DL models [37].
V-C3 Geometric Features
Geometric features can be extracted to utilize information such as length and width. During training with annotated data sets, these features can be drawn from the AIS information [32]. During operational classification, some bounding box algorithm can be used [41], as shown in Fig. 4, provided that the pixel per meter ratio is known. Length and width can be used to extract additional features [37], such as area, perimeter, aspect ratio, shape complex etc. Adding this information about geometric features can help DL models to better predict the ship classes.
Figure 4: Bounding Box of a ship target in a SAR image chip.
V-D Training Techniques
Although handcrafted features can be helpful, they require high-resolution images for effective extraction, which is often not the case, as seen in the OpenSARShip dataset. To address this challenge, researchers have explored novel deep learning techniques, further explained below.
A DL model takes input in the form of an image, learns the features, and then predicts the class of the image. However, different training strategies can be attempted to achieve promising classification performances. The techniques used in the literature for better classification are described in this section.
V-D1 Feature Fusion
A DL model trains on the features that it extracts, but when it comes to SAR ship classification, the challenge of data scarcity becomes a hurdle on the way to better results. To help the model understand the data well, several researchers propose to add hand-crafted features to the DL model [31, 43, 67, 60, 68]. Deep and handcrafted features are merged and the model classifies on this merged information, thus obtaining a better classification performance [37]. A merging strategy is schematized in Fig. 5.
Figure 5: Feature Fusion
Feature fusion is a useful technique; however, it requires manual effort, as there is no existing pipeline to extract these features automatically. Additionally, there are no standardized features that consistently enhance classification performance. It is important to note that the choice of features is resource- and time-intensive and does not always lead to better results.
V-D2 Polarization fusion
SAR images can be acquired in four transmit-receive polarizations: VV, VH, HH, and HV. While most studies in the literature have focused solely on VV or VH polarizations, it is important to note that each polarization provides distinct information about the target. Utilizing multiple polarizations (e.g., VV-VH pairs) has been shown to improve classification performance. This is because combining information from different polarizations allows the model to learn more discriminative features, leading to enhanced classification accuracy. A common approach in the literature involves creating image pairs (e.g., VV-VH) and training two separate deep learning networks simultaneously, one for each polarization [51, 40, 46, 55, 28]. As illustrated in Fig. 6, the features extracted from these networks are then merged for final classification.
Figure 6: Polarization Fusion (simplified version from [46])
Exploiting the polarization information is indeed a useful technique that provides the model with more data to learn from; however, this is not always readily available. From a cost perspective, researchers must purchase two images instead of one, significantly increasing the acquisition cost. For public datasets, polarization is available only in the OpenSARShip dataset and not in the FUSARShip dataset, limiting experimentation. Furthermore, training two models doubles the training cost and increases resource requirements.
V-D3 Metric Learning
DL models rely on feature extraction, which involves iteratively adjusting the model’s weights to optimize the extracted features. Ideally, for effective classification, the features corresponding to each class should cluster tightly in a low-dimensional space, while remaining well-separated from features of other classes. However, standard training procedures may not inherently prioritize this inter-class separation. In such cases, metric learning techniques are employed to explicitly address this challenge [30, 58, 34]. Metric learning aims to embed the extracted features into a low-dimensional space while maximizing the similarity between features belonging to the same class and minimizing the similarity between features from different classes. This is often achieved by utilizing distance metrics like cosine similarity or Euclidean distance (as illustrated in Fig. 7). Metric learning helps the model clearly distinguish between features of different classes, leading to superior performance. However, since the data is often naturally imbalanced, metric learning may be prone to overfitting. Therefore, it is crucial to balance the data beforehand or account for the imbalance when applying metric learning.
Figure 7: Metric Learning
V-D4 Multi-Scale/Pyramid Training
A DL model consists of several layers; each layer learns different scaled features, and the model makes the prediction based on the features of the last convolution layers [62, 61, 36, 33, 64, 40]. In pyramid training, see Fig. 8, the features from all the layers are directly used in decision-making by merging them using concatenation or other statistical measures (e.g. average, addition etc.).
Figure 8: Pyramid networks
Although the multiscale technique is effective at introducing features from different layers, providing the model with more data to learn from, deciding which layers to use is a critical aspect. Including all layers may introduce irrelevant information, leading the model to learn unhelpful patterns. Additionally, the method of merging features significantly impacts performance. For instance, concatenating features increases the feature size, while summing features may result in the loss of important information. Features from different scales often have varying sizes, which must be upsampled or downsampled for merging. Therefore, adopting an appropriate approach to align feature sizes is crucial for achieving better classification performance.
V-D5 Transfer-Learning and Domain Adaptation
SAR data are expensive and difficult to access, while DL models require large amounts of data. A deep neural network learns how to extract features and then applies classification to these learned features. Transfer learning can be beneficial in such scenarios. It allows leveraging other datasets to help a network learn feature extraction and then fine-tune the network with the SAR data. Transfer learning can be applied using various ways, see Fig. 9:
    Using an optical image dataset, such as ImageNet, to pre-train a network, then fine-tuning on the SAR dataset [22, 24, 44, 52].
    Utilizing open-access SAR datasets for pre-training before fine-tuning [47, 61].
    Merging optical and SAR imagery to see whether the model performs better at feature extraction, thus leading to improved overall performance [58, 45, 65, 63].
Figure 9: Transfer Learning
Transfer learning comes with the challenge of selecting the appropriate model, as numerous architectures are pretrained on large datasets. However, these models may not perform well on SAR ship classification due to the inherent imbalance in SAR data. Additionally, the depth of the networks can lead to varying classification performances. Furthermore, the nature of the data on which the model was pretrained can significantly impact its performance. Therefore, it is crucial to carefully evaluate the pretraining dataset before utilizing any pretrained network.
V-D6 Learning Rate management
One crucial component of DL is the learning rate, which helps the network to update its weights efficiently. Traditionally, the learning rate is kept fixed during the entire training. However, as a deep neural network progresses through training iterations, it gradually learns information from the data. A constant learning rate in this scenario can slow down training and potentially cause the model to miss the point of lowest error. To address this issue, a strategy was introduced by [47] that changes the learning rate based on the number of iterations. This approach reduces the learning rate as training progresses, facilitating convergence towards the optimal weights. This technique not only reduces the training time but can also improve the overall performance of deep neural networks. In tasks like SAR ship classification, where data availability is limited, such dynamic learning rate adjustments can be particularly beneficial for achieving better performances.
V-D7 Searched Networks
The architecture of a DL model is traditionally determined through the researcher’s intuition or by replicating existing architectures. However, Neural Architecture Search (NAS) offers a more objective approach to identify the optimal architecture for a specific task. NAS is a search algorithm that explores a range of possible network configurations to identify the one that performs best for the given task. This approach empowers researchers to discover the most suitable architecture for their data, potentially leading to significant advancements in maritime monitoring [50, 53]. While searched networks are an excellent option, they come with a significant computational cost. The more the layers included in the search process, the higher the computational demand. Therefore, it is crucial to carefully consider the computational cost before applying NAS.
VI Design Considerations
VI-A Model Architectures
For designing custom CNNs, researchers have proposed shallow DL architectures. These architectures typically consist of 3-5 convolutional layers followed by pooling and activation functions, ultimately connecting to 2-3 fully connected linear layers. NAS can be a valuable tool to identify the optimal architecture for a specific task. In contrast, other proposed architectures leverage established pre-trained models:
    AlexNet [79] is a pioneering deep CNN with 8 layers: 5 convolutional layers followed by 3 fully connected layers. It introduced the use of ReLU activation functions and dropout for regularization.
    VGGNet [80] features a 16-layer version, VGGNet-16, a deep CNN architecture including 13 convolutional layers and 3 fully connected layers. The convolutional layers use small 3x3 receptive fields and are stacked on top of each other to increase the depth of the network. The network also includes five max pooling layers that follow some of the convolutional layers. Its 19-layer version, VGGNet-19, includes three additional convolutional layers. Like VGGNet-16, VGGNet-19 uses small 3x3 receptive fields and includes max pooling layers.
    ResNet [81] features a 50-layer and a 34-layer versions, ResNet-50 and Resnet-34, respectively, utilizing residual learning to mitigate the vanishing gradient problem commonly encountered in deep networks. This architecture is composed of convolutional layers organized into residual blocks, each containing shortcut connections that bypass one or more layers. These shortcut connections enable the network to learn identity mappings, thus facilitating the training of much deeper networks by improving gradient flow and convergence.
    DenseNet-101, or Dense Convolutional Network [82], comprises 101 layers with a unique dense connectivity pattern. Each layer receives inputs from all preceding layers and passes its own feature maps to all subsequent layers, promoting feature reuse and mitigating the vanishing gradient problem. This dense connectivity reduces the number of parameters and encourages the network to learn more compact and efficient representations.
    InceptionNet, also known as GoogLeNet [83], introduces the inception module, which allows for multi-scale processing by applying parallel convolutional filters of different sizes ($1\times 11 × 1$, $3\times 33 × 3$ and $5\times 55 × 5$) within the same layer. This architecture comprises 22 layers, with inception modules stacked sequentially to capture diverse spatial features. This design efficiently manages computational resources while enhancing the network’s ability to learn complex and varied patterns in the input data.
These models are often adapted for the SAR image data through transfer learning approaches. Additionally, generative adversarial networks have also been explored for potential applications in SAR ship classification [45, 63].
VI-B Train-Test Split
The literature commonly uses an 80-20 split for training and testing, respectively. However, unbalanced data poses a challenge for model learning. An effective approach, as seen in the literature [55, 53], involves using the class with the fewest samples to determine the size of the training set. The remaining data are used for testing. An example of train-test ratio for class imbalanced dataset is visualized in Table IV. This approach ensures that the model learns balanced information from all classes. However, in cases of extreme imbalance, the training set becomes very limited. Therefore, it is crucial to understand the data complexity before applying such an approach.
Without considering Class Imbalance
Class
Total
100%
Train
80%
Test
20%
Cargo 1500 1200 300
Tanker 900 720 180
Fishing 600 480 120
After considering Class Imbalance
Class
Total
Train
Test
Cargo 1500 480 1020
Tanker 900 480 420
Fishing 600 480 120
TABLE IV: Numbers of ship-type occurrences for imbalanced and balanced train-test splits (example)
VI-C Hyperparameters
VI-C1 Epochs
In deep learning, an epoch is a fundamental unit of training, defined as one complete iteration over the entire training set. During each epoch, the neural network processes every sample in the dataset, enabling the model to learn and adjust its parameters—such as weights and biases—based on the error between its predictions and the actual target values. The role of epochs is critical in the iterative optimization process, as each epoch provides an opportunity for the model to refine its internal representations and progressively minimize the loss function.
The importance of epochs lies in balancing the trade-off between underfitting and overfitting; a sufficient number of epochs ensures that the model captures the underlying patterns in the data, while avoiding excessive training that might lead to overfitting, where the model performs well on the training data but poorly on new, unseen data. Properly tuning the number of epochs, often through techniques like early stopping, is essential for achieving optimal model performance and generalization. The literature suggests a common range of 50–200 epochs for DL model convergence, but the optimal number ultimately depends on the chosen learning strategy [66, 77, 58, 63, 62, 61, 60].
VI-C2 Batch Size
In deep learning, batch size refers to the number of training samples processed before updating the model’s parameters. It affects training stability and efficiency: smaller batch sizes offer more frequent updates and can improve generalization, but may introduce noise, while larger batch sizes provide more stable and accurate gradient estimates but require more memory and can slow down the training per epoch. Thus, selecting an appropriate batch size is crucial for balancing training speed, computational resources, and model performance. In the literature, batch sizes of 64, 32 and 16 are commonly used [67, 77, 64, 62, 61, 60, 49].
VI-C3 Image Size
Selection of Image size plays a crucial role in determining the network computational requirements, memory usage, and ability to capture detailed features. Larger image sizes can provide more detailed spatial information, enhancing the model’s ability to detect fine-grained patterns, but they also significantly increase computational load and memory consumption. Conversely, smaller image sizes reduce computational demands and memory usage but may result in the loss of important visual details, potentially degrading the model performance. The literature commonly utilizes image sizes of $224\times 224224 × 224$, $64\times 6464 × 64$ and $128\times 128128 × 128$ [67, 66, 64, 63].
VI-C4 Learning Rate
The learning Rate, see also Section V-D6 above, determines the step size at each iteration while moving towards a minimum of a loss function. It controls the weights of a DL network with respect to the loss gradient. A high learning rate can accelerate training by making large updates to the parameters, but risks overshooting the optimal values, leading to divergence or instability. Conversely, a low learning rate ensures more precise adjustments, promoting stable convergence, but can make the training process exceedingly slow and may get stuck in local minima. Therefore, selecting an appropriate learning rate is critical for achieving efficient and effective training. In the literature, the common learning rate is usually set in the range 0.0001 to 0.001 [67, 66, 58, 64, 63, 60]; moreover, learning rate management can be employed to train the model faster and more precisely [62, 61].
VI-D Number of Classes
Due to the imbalanced dataset problem in SAR ship data, an increase in the number of classes results in fewer samples per class. For instance, in both the public datasets considered here (FUSARShip and OpenSARShip), over 80% of the data are concentrated in just five classes on average, with the remaining data sparsely distributed among other classes. This imbalance causes classes with fewer samples to have negligible impact during training, leading the model to overfit towards the more prevalent classes. Consequently, most research focuses on 3-5 classes; although some studies have explored using more than five classes, model performance significantly deteriorates as the number of classes increases.
VII Outlook
This survey provides a comprehensive review of various techniques employed to train DL models for ship classification using SAR imagery. Through a systematic literature review, we meticulously assessed the content of 161 papers, rigorously narrowing down our selection to the most pertinent 49 studies. Their references are collected in Table II and their features are summarized in Section IV, offering an overview of their key elements and contributions to the field.
The literature reveals a broad range of DL techniques employed for SAR ship classification, as illustrated in Figure 2. These techniques include two types of architectures: CNNs, the earliest proposals, and state-of-the-art architectures such as ResNet, VGG, DenseNet etc., each tailored to leverage the unique attributes of SAR images.
Our analysis indicates that the choice of the DL technique significantly impacts classification performance. For instance, shallow CNNs have been particularly effective to learn spatial hierarchies from SAR images, as the data are scarce.
The FUSARShip and OpenSARShip datasets, along with other distinct data sources, are extensively used in the literature. These datasets are crucial for training and evaluating DL models. However, the limited availability of SAR images poses a significant challenge. To address this, data augmentation techniques have been proposed and are widely adopted. These techniques artificially increase the diversity and quantity of training data, which is essential for improving the performance and robustness of DL models. Fine-tuning pre-trained DL models on SAR-specific datasets has emerged as a successful strategy. This approach leverages the knowledge from pre-trained models and adapts it to the specific characteristics of SAR data, resulting in improved classification accuracy.
The issue of data imbalance has been addressed in the literature through two main solutions: adjusting train-test ratios and employing tailored loss functions. Additionally, some researchers propose the use of architecture search techniques like NAS for selecting the optimal DL architecture. This survey also highlights the significance of various training techniques, including feature fusion, polarization fusion, metric learning, multi-scale/pyramid learning, and learning rate management.
The findings of our survey underscore the importance of data quality and the choice of techniques in SAR ship classification. The integration of handcrafted features, effective data augmentation, and fine-tuning of pre-trained models are all critical factors contributing to the success of DL models in this domain. Building on the findings of this survey and acknowledging its limitations, several exciting avenues exist for future research in SAR ship classification using DL.
VII-A Addressing Data scarcity
The data used by researchers for DL-based SAR ship classification is seldom made public due to privacy or digital rights reasons. Future research should focus on developing a methodology to publicize the data without any such problem. Also, developing more robust data augmentation techniques can result in diverse SAR datasets. Collaborative efforts of standardizing and sharing the data can significantly benefit the research community.
VII-B Novel DL architectures
The development of DL architectures specifically tailored for SAR ship classification holds significant promise for advancing the research in this field. This includes exploring and optimizing various DL architecture components, such as loss functions, learning rates, activation functions, and even utilizing techniques like architecture search and fine-tuning. Furthermore, the continuous evolution of DL architectures presents an exciting opportunity to investigate the efficacy of novel models in SAR ship classification tasks. Recent advancements in areas such as transformers and generative models (e.g., TransGAN [84]) are particularly promising for tasks such as noise reduction and data augmentation, which can be crucial for improving the performance of DL models in SAR ship classification.
VII-C Handcrafted and Deep Features Integration
DL models have achieved remarkable success in feature learning for various tasks. However, handcrafted features, meticulously designed to capture specific characteristics of SAR images, remain a valuable tool. Further research is warranted to explore how these can be effectively integrated with DL models for SAR image classification. This combined approach has the potential to leverage the strengths of both machine learning and DL methods, potentially leading to significant performance improvements.
VII-D Standardized Performance Metrics
The field of SAR ship classification currently lacks a standardized set of performance metrics, hindering comprehensive comparisons between different studies. Utilizing metrics like F1-score can partially address this issue, but future research should focus on establishing a common set of metrics specifically tailored to the nuances of SAR ship classification tasks. This would enable more robust evaluations and facilitate meaningful comparisons between different deep learning models and training strategies.
Furthermore, inconsistencies in data utilization across studies pose another challenge to comparison. Even when researchers employ the same datasets, variations in sample selection can significantly impact the results. Moreover, among the researchers in SAR ship classification, code sharing is not common; if we want to progress further and faster, code sharing should be adopted as a practice. Standardized metrics and data utilization practices are crucial for achieving truly comparable research, ultimately accelerating the advancements in the field.
VII-E Interpretability of DL models
DL models, while demonstrably powerful, are often characterized as “black boxes” due to their complex internal workings. This lack of interpretability can hinder our understanding of the rationale behind their classification decisions. Fortunately, various interpretability techniques, such as Gradient-weighted Class Activation Mapping (Grad-CAM, [85]), SHapley Additive exPlanations (SHAP, [86]) and Local Interpretable Model-Agnostic Explanations (LIME, [87]) can be employed to shed light on the internal decision-making processes of these models.
While interpretability has not always been a central focus in SAR ship classification research, it presents a valuable avenue for future exploration. By leveraging interpretability techniques, researchers can gain deeper insights into the features most critical for the model’s decisions. This enhanced understanding can lead to more informed analysis, improved trust in model predictions, and potentially even the identification of potential biases within the model.
VII-F Interdisciplinary Collaborations
Significant advancements in SAR ship classification using DL can be achieved by fostering interdisciplinary collaborations between researchers in remote sensing, computer vision, and maritime studies. This collaborative approach has the potential to not only generate innovative solutions for the challenges faced in SAR ship classification, but also contribute to environmental protection efforts by facilitating the detection of illegal maritime activities.
By focusing on these directions, future research can build on our findings and continue to advance the field of SAR ship classification using deep learning, ultimately improving the accuracy and efficiency of maritime surveillance and related applications.
VIII Conclusion
This survey has systematically reviewed the diverse techniques employed to train deep learning models for SAR ship classification. We identify critical trends and challenges in this domain, emphasizing the importance of integrating handcrafted features, utilizing public datasets, data augmentation, fine-tuning, explainability, and interdisciplinary collaborations for enhancing DL model performance.
We established a survey methodology with clearly defined research questions. Our literature search yielded a comprehensive dataset of relevant papers, which were then categorized into a first-of-its-kind taxonomy encompassing deep learning techniques, handcrafted feature use, SAR attribute utilization, and the impact of fine-tuning.
Section IV provides detailed explanations of the included papers, categorized into the four sections (discussing DL techniques, impact of handcrafted features, use of SAR properties and fine-tuning). Section V presents the methodologies employed in SAR ship classification. Finally, Section VII outlines potential avenues for addressing current challenges, such as data scarcity, novel DL architectures, interpretability, performance metrics, feature integration, and interdisciplinary collaborations.
Advancements in DL techniques for SAR ship classification hold significant promise for improving maritime surveillance and related applications. Continued exploration and innovation in this field can lead to more accurate and efficient ship classification systems, ultimately enhancing the capabilities of maritime operations.
Acknowledgments
This work was supported by National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.4 - Call for tender No. 3138 of 16 December 2021, rectified by Decree n.3175 of 18 December 2021 of Italian Ministry of University and Research funded by the European Union – NextGenerationEU. Award Number: Project code CN_00000033, Concession Decree No. 1034 of 17 June 2022 adopted by the Italian Ministry of University and Research, CUP D33C22000960007, Project title “National Biodiversity Future Center - NBFC”.